# Common Envelope Evolution of Ultralight Boson Clouds

### v1

@@ I Introduction
Ultralight bosons can arise in many fundamental theories [1], and are proposed as possible dark matter candidates [2, 3, 4, 5, 6, 7]. If the Compton wavelength of a boson field is comparable to the size of a rapidly rotating black hole, the field can experience superradiant growth, forming a cloud around the black hole [8, 9]. The resulting system, often referred to as a gravitational atom due to its structural similarity to a hydrogen atom, provides a promising laboratory for probing ultralight bosons [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21].
@@
Clouds of ultralight bosons can also form around black holes in a binary system. The evolution of such clouds [22, 23, 24, 25, 26, 27, 28, 29, 30] as well as their signatures on orbital dynamics [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44] have been extensively studied in the literature. In particular, a cloud in a binary is expected to be bounded by its host black hole, when the orbital separation is much larger compared to the cloud, or if the host black hole is much heavier than the companion black hole; cf. Fig. 1. In these cases, the gravitational effects of the companion black hole can be treated perturbatively, and a cloud that is initially in one of the bound states can transit to other bound states [22, 23] and even ionization [28] under the tidal perturbation of the companion black hole. On the other hand, for comparable mass ratio binaries, it is shown in Refs. [44, 29] that the cloud bounded by one of the binary black holes can start transferring to the region around another black hole, as the orbital separation approaches the size of the cloud. Eventually, one may expect that the cloud (or clouds if both black holes are dressed with their own clouds, which is very likely for comparable mass ratio binaries) will form a common envelope, surrounding the whole binary. In this case, the system may no longer be treated as a perturbed gravitational atom but is more like a rotating gravitational molecule [25].
@@
In this work, we investigate the formation and evolution of the common envelope state of an ultralight boson cloud in comparable mass-ratio binaries. Focusing on cloud evolution in the late stage, i.e., when the orbital separation is comparable to the cloud size, we describe the cloud evolution with the molecular eigenstates, and investigate the possible level transitions between the eigenstates during orbital decay and the potential effects of cloud on orbital dynamics.
@@
This paper is organized as follows: We first introduce gravitational molecules in Sec. II, highlighting their relation to the gravitational atoms. Then we set up the framework for analyzing the evolution of gravitational molecules in circular orbits in Sec. III and in elliptical orbits in Sec. IV. Finally, we discuss the common envelope evolution of superradiant clouds in Sec. V and its signatures on orbital dynamics in Sec. VI. Sec. VII devotes to conclusion and discussion. We will take (,+,+,+)(-,+,+,+) metric convention and set $\hbar=c=1$. We use $\varphi$ for atomic states and $\psi$ for molecular states.
@@
Figure 1: Illustration of cloud common envelope formation. For comparable mass-ratio binaries, a cloud is expected to be bounded by its own host black holes when the orbital separation is much larger comparing to the cloud. As the orbit decays, the clouds may undergo mass transfer and eventually form a common envelope surrounding the binary.
@@ II Gravitational Molecule
In this section, we shall introduce gravitational molecule, which is proposed in Ref. [25]. We shall consider a real scalar field $\Phi$ of mass $\mu$ in a black hole binary spacetime, the dynamics of which is governed by Klein-Gordon equation,
@@
is the Newtonian potential, and $M_{1,2}$ and $\mathbf{r_{1,2}}$ denote the mass and position of each black hole. In the non-relativistic limit, we take the ansatz
@@
where $\Psi$ is a complex field that varies on a time scale much longer than $\mu^{-1}$. Substituting Eq. (4) into Eq. (1) and keeping only the leading terms in the non-relativistic and weak field limits, the Klein-Gordon equation (1) reduces to a Schrödinger equation,
@@
The Newtonian potential $\Phi_{N}$ is time dependent. Nevertheless, we can remove the time dependence by working in a co-rotating frame in the case of circular orbits: We first introduce the lab frame $(t,r,\theta,\phi)$ such that the origin is centered at the center of mass of the binary, and the binary orbit lies in the equatorial plane, and then define the co-rotating frame $(t,\bar{r},\bar{\theta},\bar{\phi})$ with
@@
where $\Omega$ is the orbital frequency. In the co-rotating frame, Eq. (5) is rewritten as
@@
As pointed in Ref. [25], in the absence of $H^{\prime}$, Eq. (7) reduces to
@@
which resembles the Schrödinger equation of the electron in a single electron heteronuclear diatomic molecule. While the eigenstates of Eq. (10) can be obtained numerically for a general mass ratio $q=M_{2}/M_{1}$, we will focus on equal mass binaries, i.e., $q=1$, in most of this work. As we will show later, the asymptotical properties of the eigenstates can be derived analytically for $q=1$, making the equal mass binaries a good benchmark.
When $q=1$, $\bar{\Psi}$ resembles the wavefunction of the electron in a hydrogen molecule ion with a “fine structure constant” $\alpha\equiv GM\mu$ and a “Bohr radius” $r_{B}=GM/\alpha^{2}$. The bound states are known to be discrete in spectrum with each eigenstate labeled by three “quantum numbers”,
@@
The energy and the wavefunction of the eigenstates depend on the separation between two black holes $\left|{\bar{\mathbf{r}}}_{2}-{\bar{\mathbf{r}}}_{1}\right|\equiv a$. As $a$ approaches zero, the eigenstates approach the eigenstates of the hydrogen atom $\bar{\psi}_{n\ell m}^{(0)}$. It is just that the Bohr radius of the approached hydrogen atom is $r_{B}/4$, not $r_{B}$. On the other hand, in the limit of $a\gg r_{B}$, the clouds should be bounded around the two black holes separately. In this case, a molecular eigenstate $\bar{\psi}^{(\infty)}_{n\ell m}$ can be approximated by the superposition of two identical atomic eigenstates,
@@
where $\bar{\varphi}_{n_{s}km}$ is the eigenstate of a gravitational atom. Note that the wavefunction $\bar{\varphi}_{n_{s}km}$ is obtained in the parabolic coordinates centered on the $i$-th black hole, i.e., $\left\{\rho_{i},\sigma_{i},\chi\right\}$, and is labeled by $n_{s}$, $k$ and $m$, which are the principal, parabolic and azimuthal quantum numbers, respectively (see Ref. [45] for details).
Figure 2: Energy levels of the molecular eigenstates. The vertical and horizontal axes show the dimensionless energy $\tilde{\epsilon}\equiv\epsilon_{n\ell m}/\mu\alpha^{2}$ and dimensionless orbital semi-major axis $\tilde{a}\equiv a/r_{B}$. The upper and the lower plot show the energy levels with $q=1$ and $q=2/3$ respectively. For $q=1$, the system is symmetric under reflection, which breaks when $q\neq 1$. As a result, each energy level in the upper plot splits into two levels in the lower plot.
Although the molecular eigenstates and the atomic eigenstates in Eq. (12) are labeled with different sets of quantum numbers, they, however, can be related by conserving the number of nodal surfaces of the eigenfunctions [46], [^1] The relation used in this paper is different from the one in Ref. [46]. When $|m|$ is even, the superposition of two well-separated atomic states, i.e., $\bar{\varphi}_{1,n_{s}km}\pm\bar{\varphi}_{2,n_{s}km}$ can introduce or eliminate an additional node on the nodal surface where $\zeta=0$, which modifies the relation given by Eq.(19) and Eq.(20) of Ref. [46].
@@
The upper plot in Fig. 2 shows the energy of some molecular eigenstates, where the solid lines correspond to the numerical results obtained with the continued fraction method [47] and the shooting method [48] (see App. A for details). The numerical results show how the spectrum splits as the orbital separation varies and confirm the relations (13) between the quantum numbers. To avoid confusion, we shall label a molecular eigenstate with $n$, $\ell$ and $m$ hereafter, which are also the quantum numbers of the approached atomic eigenstate in the small separation limit. We also show the energy levels with mass ratio $q=2/3$, in which case the reflection symmetry breaks, and each $\epsilon_{n\ell m}$ in the case of $q=1$ splits into two levels once $a$ deviates from 0 as shown in the plot.
III Could Evolution in Circular Orbits
@@
where $i$ is a short notation for the quantum numbers of the molecular eigenstate. [^2] In principle, one cloud also include unbound states, transitions which states could become efficient when $a\sim r_{B}$. In this paper, we shall focus on interactions between bound states, and will investigate transitions to unbound states in future study. In this case, the cloud evolution is fully described by $c_{i}(t)$. However, in practice, the orbital separation also evolves with time. If the orbit evolution is dominated by gravitational wave radiation, the orbital frequency changes on a time scale
@@
The first rule arises from the periodicity of $\chi$, while the second rule is due to the parity of the eigenstates given $m_{i}-m_{j}=\pm 1$. In the small separation limit, the coupling strength approaches to
@@
where $d^{\ell}_{mm^{\prime}}\left(\frac{\pi}{2}\right)$ is an element of the Wigner-D matrix. In the large separation limit, the wavefunction is nontrivial only at $\zeta^{2}\sim 1-2r_{B}/a$ and $\xi^{2}\sim 1+2r_{B}/a$. As a result, $\partial_{\bar{\phi}}$ is dominated by the second term in Eq. (17), and thus $\eta_{ij}/\Omega$ scales with $a/r_{B}$ at large separation limit. In Fig. 3, we show the coupling strengths obtained with numerical integrations. We find that the coupling strengths approach to zero in small separation limit if the two modes have different $n$ or $\ell$, and all scale with $a$ in large separation limit, which is consistent with the discussion above.
@@
Estimating energy difference between the molecular states with $\sim\alpha^{3}/GM$, one can find $\left|\eta_{ij}/\epsilon_{ij}\right|\sim(r_{B}/a)^{1/2}$, which means each eigenstate evolves adiabatically when $a\gg r_{B}$ even in the presence of the $H^{\prime}$. However, nonadiabatic transitions could occur when the energy levels of two coupled states cross each other at a certain orbital separation $a_{\times}$ as the orbital evolves. This process is known as the Landau-Zener transition. The transition rate can be estimated as $P_{ij}^{\times}=e^{-2\pi Z_{ij}^{\times}}$ with [49, 50]
@@
IV Could Evolution in Eccentric orbits
@@
where $\phi_{*}(t)$ is the true anomaly of the orbit and $a$ is the semi-major axis that characterizes the orbit separation. For $e\ll 1$, we can expand $H$ in $e$. By doing so, we have
@@
being the averaged orbital frequency. Taking ansatz (14), the evolution of cloud in eccentric orbits is governed by
@@
For equal mass binaries, the first term in Eq. (26) is proportional to $\eta_{ij}$, which has been discussed in Sec. III with values shown in Fig. 3 for some couplings. The second term in Eq. (26), on the other hand, turns out to be zero when $q=1$. This is because the selection rule with $m_{i}=m_{j}\pm 1$ requires that the coupled two states have to be of one symmetric and one anti-symetric with respect to the plane $\phi=\pm\pi/2$, in which case the gravitational effects of the two black holes cancel in the integral.
@@
According to Eq. (25), a non-zero eccentricity could also lead to resonant transitions between two molecular eigenstates. Those transitions occur at $\bar{\Omega}=\pm\epsilon_{ij}$ if $\eta^{\prime}_{ij}\neq 0$. Near the resonant orbit, that is, orbit with $\bar{\Omega}\simeq\pm\epsilon_{ij}$, one can neglect the highly oscillating terms proportional to $e^{i\epsilon_{ij}t}$ and $e^{i\epsilon_{ij}t\mp i\bar{\Omega}t}$. The transition is then Landau-Zener type, and the transition rate can be estimated by $P_{ij}^{\,e}=e^{-2\pi Z_{ij}^{e}}$, where

### v2

@@ I Introduction
Ultralight bosons can arise in many fundamental theories [1] and are proposed as possible dark matter candidates [2, 3, 4, 5, 6, 7]. If the Compton wavelength of a boson field is comparable to the size of a rapidly rotating black hole, the field can experience superradiant growth, forming a cloud around the black hole [8, 9, 10, 11, 12, 13]. The resulting system, often referred to as a gravitational atom due to its structural similarity to a hydrogen atom, provides a promising laboratory for probing ultralight bosons [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25].
@@
Clouds of ultralight bosons can also form around black holes in a binary system. The evolution of such clouds [26, 27, 28, 29, 30, 31, 32, 33, 34, 35] as well as their signatures on orbital dynamics [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53] have been extensively studied in the literature. In particular, a cloud in a binary is expected to be bounded by its host black hole, when the orbital separation is much larger compared to the cloud, or if the host black hole is much heavier than the companion black hole; cf. Fig. 1. In these cases, the gravitational effects of the companion black hole can be treated perturbatively, and a cloud that is initially in one of the bound states can transit to other bound states [26, 27] and even ionization [32] under the tidal perturbation of the companion black hole. On the other hand, for comparable mass-ratio binaries, it is shown in Refs. [49, 33] that the cloud bounded by one of the binary black holes can start transferring to the region around another black hole, as the orbital separation approaches the size of the cloud. Eventually, one may expect that the cloud (or clouds if both black holes are dressed with their own clouds, which is very likely for comparable mass-ratio binaries) will form a common envelope, surrounding the whole binary. In this case, the system may no longer be treated as a perturbed gravitational atom but is more like a rotating gravitational molecule [29].
@@
In this work, we investigate the formation and evolution of the common envelope state of an ultralight boson cloud in comparable mass-ratio binaries. Focusing on cloud evolution in the late stage, i.e., when the orbital separation is comparable to the cloud size, we describe the cloud evolution with the molecular eigenstates and investigate the possible level transitions between the eigenstates during orbital decay and the potential effects of the cloud on orbital dynamics.
@@
This paper is organized as follows: We first introduce gravitational molecules in Sec. II, highlighting their relation to the gravitational atoms. Then we set up the framework for analyzing the evolution of gravitational molecules in circular orbits in Sec. III and in elliptical orbits in Sec. IV. Finally, we discuss the common envelope evolution of superradiant clouds in Sec. V and its signatures on orbital dynamics in Sec. VI. Section  VII devotes to conclusion and discussion. We will take the (,+,+,+)(-,+,+,+) metric convention and set $\hbar=c=1$. We use $\varphi$ for atomic states and $\psi$ for molecular states.
@@
Figure 1: Cloud common envelope formation. For comparable mass-ratio binaries, a cloud is expected to be bounded by its own host black holes when the orbital separation is much larger compared to the cloud. As the orbit decays, the clouds may undergo mass transfer and eventually form a common envelope surrounding the binary.
@@ II Gravitational Molecule
In this section, we shall introduce the gravitational molecule, which is proposed in Ref. [29]. We shall consider a real scalar field $\Phi$ of mass $\mu$ in a black hole binary spacetime, the dynamics of which is governed by Klein-Gordon equation
@@
is the Newtonian potential, and $M_{1,2}$ and $\mathbf{r_{1,2}}$ denote the mass and position of each black hole. In the nonrelativistic limit, we take the ansatz
@@
where $\Psi$ is a complex field that varies on a timescale much longer than $\mu^{-1}$. Substituting Eq. (4) into Eq. (1) and keeping only the leading terms in the nonrelativistic and weak field limits, the Klein-Gordon equation (1) reduces to a Schrödinger equation,
@@
The Newtonian potential $\Phi_{N}$ is time dependent. Nevertheless, we can remove the time dependence by working in a corotating frame in the case of circular orbits: We first introduce the lab frame $(t,r,\theta,\phi)$ such that the origin is centered at the center of mass of the binary, and the binary orbit lies in the equatorial plane, and then we define the corotating frame $(t,\bar{r},\bar{\theta},\bar{\phi})$ with
@@
where $\Omega$ is the orbital frequency. In the corotating frame, Eq. (5) is rewritten as
@@
As pointed out in Ref. [29], in the absence of $H^{\prime}$, Eq. (7) reduces to
@@
which resembles the Schrödinger equation of the electron in a single electron heteronuclear diatomic molecule. While the eigenstates of Eq. (10) can be obtained numerically for a general mass ratio $q=M_{2}/M_{1}$, we will focus on equal-mass binaries ($M_{1}=M_{2}=M$ and $q=1$), in most of this work. As we will show later, the asymptotical properties of the eigenstates can be derived analytically for $q=1$, making the equal-mass binaries a good benchmark.
When $q=1$, $\bar{\Psi}$ resembles the wave function of the electron in a hydrogen molecule ion with a “fine structure constant” $\alpha\equiv GM\mu$ and a “Bohr radius” $r_{B}=GM/\alpha^{2}$. The bound states are known to be discrete in spectrum with each eigenstate labeled by three “quantum numbers,”
@@
The energy and the wave function of the eigenstates depend on the separation between two black holes $\left|{\bar{\mathbf{r}}}_{2}-{\bar{\mathbf{r}}}_{1}\right|\equiv a$. As $a$ approaches zero, the eigenstates approach the eigenstates of the hydrogen atom $\bar{\psi}_{n\ell m}^{(0)}$. It is just that the Bohr radius of the approached hydrogen atom is $r_{B}/4$, not $r_{B}$. On the other hand, in the limit of $a\gg r_{B}$, the clouds should be bounded around the two black holes separately. As a result, the Hilbert space of the gravitational molecule is the direct sum of the two gravitational atoms’ Hilbert spaces. In this case, a molecular eigenstate $\bar{\psi}^{(\infty)}_{n\ell m}$ can be approximated by the superposition of two identical atomic eigenstates,
@@
where $\bar{\varphi}_{n_{s}km}$ is the eigenstate of a gravitational atom. Note that the wave function $\bar{\varphi}_{n_{s}km}$ is obtained in the parabolic coordinates centered on the $i$th black hole, i.e., $\left\{\rho_{i},\sigma_{i},\chi\right\}$, and is labeled by $n_{s}$, $k$, and $m$, which are the principal, parabolic and azimuthal quantum numbers, respectively (see Ref. [54] for details).
Figure 2: Energy levels of the molecular eigenstates. The vertical and horizontal axes show the dimensionless energy $\tilde{\epsilon}\equiv\epsilon_{n\ell m}/\mu\alpha^{2}$ and dimensionless orbital semimajor axis $\tilde{a}\equiv a/r_{B}$. The upper and the lower plots show the energy levels with $q=1$ and $q=2/3$ respectively. For $q=1$, the system is symmetric under reflection, which breaks when $q\neq 1$. As a result, each energy level in the upper plot splits into two levels in the lower plot.
Although the molecular eigenstates and the atomic eigenstates in Eq. (12) are labeled with different sets of quantum numbers, they, however, can be related by conserving the number of nodal surfaces of the eigenfunctions [55], [^1] The relation used in this paper is different from the one in Ref. [55]. When $|m|$ is even, the superposition of two well-separated atomic states, i.e., $\bar{\varphi}_{1,n_{s}km}\pm\bar{\varphi}_{2,n_{s}km}$ can introduce or eliminate an additional node on the nodal surface where $\zeta=0$, which modifies the relation given by Eqs.(19) and (20) of Ref. [55].
@@
The upper plot in Fig. 2 shows the energy of some molecular eigenstates, where the solid lines correspond to the numerical results obtained with the continued fraction method [56] and the shooting method [57] (see Appendix  A for details). The numerical results show how the spectrum splits as the orbital separation varies and confirm the relations (13) between the quantum numbers. To avoid confusion, we shall label a molecular eigenstate with $n$, $\ell$, and $m$ hereafter, which are also the quantum numbers of the approached atomic eigenstate in the small separation limit.
In the case of $q\neq 1$, the molecular energy spectrum in the limit of $a\rightarrow\infty$ is still the union of the two energy spectra of two gravitational atoms, but the two atomic energy spectra are no longer identical. The break in the reflection symmetry leads to the energy level split for a given set of ($n$, $\ell$, $m$). To understand the level split, we first recall the case of $q=1$. In the limit of $a\rightarrow\infty$, the atomic states $\varphi_{1,n\ell m}$ and $\varphi_{2,n\ell m}$ are also molecular eigenstates, and they are degenerate. (Note that they are no longer molecular eigenstates for a finite $a$.) As $a\rightarrow 0$, these two states correspond to the same molecular eigenstate. Now we consider the case of $q\neq 1$. The two atomic states are still molecular eigenstates as $a\rightarrow\infty$, and correspond to the same molecular state as $a\rightarrow 0$, but they are generally not degenerate. Therefore, the energy level denoted by a certain set of ($n$, $\ell$, $m$), the quantum numbers in the small separation limit, should split into two levels and $a$ deviates from 0. In the lower plot in Fig. 2, we show the energy levels with mass ratio $q=2/3$, in which case the reflection symmetry breaks, and each $\epsilon_{n\ell m}$ in the case of $q=1$ splits into two levels once $a$ deviates from 0.
III Cloud Evolution in Circular Orbits
@@
where $i$ is a short notation for the quantum numbers of the molecular eigenstate. [^2] In principle, one cloud also includes unbound states, transitions where states could become efficient when $a\sim r_{B}$. In this paper, we shall focus on interactions between bound states, and we will investigate transitions to unbound states in future study. In this case, the cloud evolution is fully described by $c_{i}(t)$. However, in practice, the orbital separation also evolves with time. If the orbit evolution is dominated by gravitational wave radiation, the orbital frequency changes on a timescale
@@
The first rule arises from the periodicity of $\chi$, while the second rule is due to the parity of the eigenstates given $m_{i}-m_{j}=\pm 1$. In the small separation limit, the coupling strength approaches
@@
where $d^{\ell}_{mm^{\prime}}\left(\frac{\pi}{2}\right)$ is an element of the Wigner D-matrix. In the large separation limit, the wave function is nontrivial only at $\zeta^{2}\sim 1-2r_{B}/a$ and $\xi^{2}\sim 1+2r_{B}/a$. As a result, $\partial_{\bar{\phi}}$ is dominated by the second term in Eq. (17), and thus $\eta_{ij}/\Omega$ scales with $a/r_{B}$ at a large separation limit. In Fig. 3, we show the coupling strengths obtained with numerical integrations. We find that the coupling strengths approach zero in a small separation limit if the two modes have different $n$ or $\ell$, and all scale with $a$ in a large separation limit, which is consistent with the discussion above.
@@
Estimating the energy difference between the molecular states with $\sim\alpha^{3}/GM$, one can find $\left|\eta_{ij}/\epsilon_{ij}\right|\sim(r_{B}/a)^{1/2}$, which means each eigenstate evolves adiabatically when $a\gg r_{B}$ even in the presence of $H^{\prime}$. However, nonadiabatic transitions could occur when the energy levels of two coupled states cross each other at a certain orbital separation $a_{\times}$ as the orbital evolves. This process is known as the Landau-Zener transition. The transition rate can be estimated as $P_{ij}^{\times}=e^{-2\pi Z_{ij}^{\times}}$ with [58, 59]
@@
IV Cloud Evolution in Eccentric orbits
@@
where $\phi_{*}(t)$ is the true anomaly of the orbit and $a$ is the semimajor axis that characterizes the orbit separation. For $e\ll 1$, we can expand $H$ in $e$. By doing so, we have
@@
being the averaged orbital frequency. Taking ansatz (14), the evolution of the cloud in eccentric orbits is governed by
@@
For equal-mass binaries, the first term in Eq. (26) is proportional to $\eta_{ij}$, which has been discussed in Sec. III with values shown in Fig. 3 for some couplings. The second term in Eq. (26), on the other hand, turns out to be zero when $q=1$. This is because the selection rule with $m_{i}=m_{j}\pm 1$ requires that the coupled two states have to be of one symmetric and one antisymetric with respect to the plane $\phi=\pm\pi/2$, in which case the gravitational effects of the two black holes cancel in the integral.
@@
According to Eq. (25), a nonzero eccentricity could also lead to resonant transitions between two molecular eigenstates. Those transitions occur at $\bar{\Omega}=\pm\epsilon_{ij}$ if $\eta^{\prime}_{ij}\neq 0$. Near the resonant orbit, that is, orbit with $\bar{\Omega}\simeq\pm\epsilon_{ij}$, one can neglect the highly oscillating terms proportional to $e^{i\epsilon_{ij}t}$ and $e^{i\epsilon_{ij}t\mp i\bar{\Omega}t}$. The transition is then Landau-Zener type, and the transition rate can be estimated by $P_{ij}^{\,e}=e^{-2\pi Z_{ij}^{e}}$, where